A girl sits at a distance of 12.0 m from the axis of merry-go-round in an amusement park. If the merry-go-round makes 3 rev


in 12 Os find the (a) tangential speed and (b) centripetal acceleration of the child.

1
Expert's answer
2022-04-06T18:36:41-0400

(a) tangential speed


"v=\frac{2\pi R}{T}=\frac{2\pi*12.0}{4.00}=18.8\:\rm m/s"

(b) centripetal acceleration


"a=\frac{v^2}{R}=\frac{18.8^2}{12.0}=29.6\:\rm m/s^2"